OpenStudy (anonymous):

f we multiply all design dimensions of an object by a scaling factor f, its volume and mass will be multiplied by f^3. Part A: By what factor will its moment of inertia be multiplied?

OpenStudy (fellowroot):

Someone jump in if I'm wrong, but doesn't it depend on the type of object you are talking about. A thin hoop for instance has a different moment of inertia than a sphere.

OpenStudy (anonymous):

Let's see from the definition of moment of inertia, \(I=\int r^2 dm\) (or just use \(I=mr^2 \) for this purpose). If we multiply r by f, and m by \(f^3\), \(I'=fm(f^3r)^2=f^7mr^2\). The new moment of inertia is f^7 times the initial.

OpenStudy (anonymous):

Sorry, I'=f^3m(fr)^2=f^5 mr^2 \(I'=f^5I\)